Tohban Report 2015-05-20

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 13 May 2015
End Date: 20 May 2015
Tohban: Hazel Bain
Tohban email:
Next Tohban: Milo Buitrago-Casas
List all reports


Solar Activity

Solar activity this week was low this week. The majority of the C class flares this week were produced by ARs 12339 and 12345, which have now rotated off the disk. There are currently ten active regions on the disk. Solar activity is currently low, with a GOES background flux at B class level.

How many GOES flares occurred?

 Flares above B, C, M, X class were      9    24     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      3    14     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      1     8     0     0

There were RHESSI flares/GOES flares 167 / 33 over the time range 13-May-15 20-May-15

Memory Management

The SSR fill level reached a maximum of ~50% and is not currently emptying at the end of the daily pass set (currently ~29% at the end of the passes).

Spacecraft Status

The cold plate 1 and 2 are at about 132.2 K and 130.7 K respectively. There was ~0.5 K step decrease in the cold tip temperature on the 16th of May.

Accelerations remain low at 11.5mg.

Data Gaps

No data gaps this week

Detector issues

- The front fast threshold on detector 9 was raised from 0x30 to 0x40.

2015-133-19:09:52 /IDPLOAD VALUE=0x40  ; was 0x30

- The Front slow threshold on detector 8 was increased from 0x0C to 0x15.

2015-135-21:36:25 /IDPLOAD VALUE=0x15  ; was 0x0C

- The front fast rates on D3, D4 and D6, and the rear fast rates on D1, D5, D6 and D7 began increasingly significantly. The trend began around the 7th of May but has become more pronounced during the last week. The trend appears to follow the rising cold tip temperature trend, but is increasing faster than in previous months and may point to another source of noise on the spacecraft. The cold tip temperature has reached it's current peak and will start to decrease again due to the changing day length. Therefore the decision was made to hold of on increasing the fast rate thresholds on all detectors and observer the current trend for a couple more days.

Other notes

Spacecraft Management

- At the start of the week the thin attenuator was stuck in due to the livetime of detectors 1 and 7 dropping below 99.2% consistently. The HV on detector 1 and 7 was dropped by 20 steps each and the thin attenuator came out at 19:02:28 UT on the 13th of May.

2015-133-19:02:13 /ihvdac detector=1, voltage=186
2015-133-19:02:24 /ihvdac detector=7, voltage=211

- At 07:30 UTC on the 14th, the thin attenuator returned to being stuck in despite the change in HV. At 20:16 UT on the 14th of May we changed the threshold used for pulling the thin shutter out, more specifically going from attenuator state 1 to 0. The threshold was changed from 0.8% to 1.2% dead time, the step size is 0.4%. Once this commanding was done the thin shutter came out. Since the change solar activity has been low and it still remains to be seen how the attenuator will react during a flare using this new setting. For now it is out and we will keep a close eye on it.

- The decimation was changed from active to normal_vigorous.

2015-133-19:08:22 start idpu_dec_normal_vigorous

Decimation Normal/Vigorous
Night time data (fronts) No nighttime events
Night time data (rears) Taking nighttime events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools